SunnyLandscape04:

How do I solve??? 2x+5+8x=22

ThisGuyHandsome:

I got this one! It will be step by step though.

supie3:

first you have to combine the like terms which are 2x and 8x

ThisGuyHandsome:

๐™๐™๐™ž๐™จ ๐™š๐™ฆ๐™ช๐™–๐™ฉ๐™ž๐™ค๐™ฃ ๐™ž๐™จ ๐™จ๐™ค๐™ก๐™ซ๐™š๐™™ ๐™—๐™ฎ ๐™ช๐™จ๐™ž๐™ฃ๐™œ ๐™ฉ๐™๐™š ๐™›๐™ช๐™ฃ๐™˜๐™ฉ๐™ž๐™ค๐™ฃ๐™จ ๐™ค๐™› ๐™˜๐™ค๐™ข๐™—๐™ž๐™ฃ๐™ž๐™ฃ๐™œ ๐™ก๐™ž๐™ ๐™š ๐™ฉ๐™š๐™ง๐™ข๐™จ. \[10x+5=22\] ๐™๐™๐™š๐™ฃ ๐™จ๐™ž๐™ฃ๐™˜๐™š ๐™ž๐™ฉโœ๐™จ ๐™ฅ๐™ค๐™จ๐™ž๐™ฉ๐™ž๐™ซ๐™š, ๐™ฎ๐™ค๐™ช ๐™จ๐™ช๐™—๐™ฉ๐™ง๐™–๐™˜๐™ฉ ๐™ฉ๐™๐™š ๐™›๐™ž๐™ซ๐™š ๐™›๐™ง๐™ค๐™ข ๐™ฉ๐™๐™š ๐™ฉ๐™ฌ๐™š๐™ฃ๐™ฉ๐™ฎ-๐™ฉ๐™ฌ๐™ค. \[10x=17\] ๐™๐™ž๐™ฃ๐™–๐™ก๐™ก๐™ฎ, ๐™™๐™ž๐™ซ๐™ž๐™™๐™š ๐™จ๐™š๐™ซ๐™š๐™ฃ๐™ฉ๐™š๐™š๐™ฃ ๐™—๐™ฎ ๐™ฉ๐™š๐™ฃ, ๐™ฌ๐™๐™ž๐™˜๐™ ๐™จ๐™๐™ค๐™ช๐™ก๐™™ ๐™œ๐™ž๐™ซ๐™š ๐™ฎ๐™ค๐™ช ๐™ฉ๐™๐™š ๐™›๐™ž๐™ฃ๐™–๐™ก ๐™–๐™ฃ๐™จ๐™ฌ๐™š๐™ง ๐™ค๐™›; \[X=1.7\] ๐™„ ๐™๐™ค๐™ฅ๐™š ๐™ฉ๐™๐™ž๐™จ ๐™œ๐™ž๐™ซ๐™š๐™ฃ ๐™ž๐™ฃ๐™›๐™ค๐™ง๐™ข๐™–๐™ฉ๐™ž๐™ค๐™ฃ ๐™๐™š๐™ก๐™ฅ๐™จ; ๐˜พ๐™ค๐™ฃ๐™ฉ๐™–๐™˜๐™ฉ ๐™ข๐™š ๐™ž๐™› ๐™ฃ๐™š๐™š๐™™๐™š๐™™. @๐™๐™๐™ž๐™จ๐™‚๐™ช๐™ฎ๐™ƒ๐™–๐™ฃ๐™™๐™จ๐™ค๐™ข๐™š

supie3:

the like terms are 2x and 8x so we would move 2x and 8x next to each other so that they can be combined so, we would have 2x+8x+5=22, and 2x+8x=10x so now we have 10x+5=22 then we change the plus to a minus in `+5` so that itโ€™s not โ€“5 and subtract 5 from both sides of the equation 22-5=17 so we would have 10x-17 then divide both sides by 10 \(\cancel{\frac{10x}{10}}-\frac{17}{10}\) (10x/10 cancels) then all thatโ€™s left is 17/10, so x=17/10

supie3:

@thisguyhandsome wrote:
i literally just did that work
you barely showed any work you didnt even show combining like terms or wtv